A manufacturing process on the critical path of a project is expected to get easier as we go along and the project manager wants to plan accurately for when he expects the task to be complete.

How many cycles are anticipated before a new operator reaches a standard production time of 50 minutes if the first unit required 101 minutes and the learning curve rate is 0.77?
Answer #1

Tn = T1 * nb

T1 = 101 minutes

b = log (r) / log 2

b = −0.37707

Tn = 50

So,

50 = 101 * n−0.37707

Solving, we get

n=6.453573

So, we can anticipate atleast 7 cycles before a new operator reaches the given condition.

PLEASE LEAVE A LIKE. IT REALLY HELPS ME A LOT. THANK YOU!!!